OpenStudy (anonymous):

Find two numbers that have a product of -40 and a sum of 3.

OpenStudy (anonymous):

8,- 5

OpenStudy (lgbasallote):

let x ad y be the numbers x + y = 3 xy = 40 let us use substitution since x + y = 3 x = 3 - y let us now substitute that to xy = 40 xy = 40 (3-y)y = 40 3y - y^2 = 40 put everything on the right side y^2 - 3y + 40 = 0 you can factor this out into (y+5)(y-8)

mathslover (mathslover):

xy=-40 x+y=3 (x+y)^2=9 x^2+y^2+2xy=9 x^2+y^2+2(-40)=0 x^2+y^2-80=0 x^2+y^2=89 (x-y)^2 = x^2+y^2-2xy = 89 - 2(-40) = 89+80 = 169 (x-y)=+13 or -13 x-y = 13 x + y = 3 2x = 16 -2y=10 x=8 y=-5

OpenStudy (anonymous):

thanks for simplifying that it really helped u guys :)

OpenStudy (anonymous):

Or you can simply write out the factors of -40: -10, 4 10,-4 8, -5 -8, 5 20, -2 -20, 2 and then testing which add up to 3

OpenStudy (anonymous):

is there anyway i can give everyone a medal?

mathslover (mathslover):

yes tell us to give a medal to which person i will give to other one..

OpenStudy (anonymous):

okay well can @mathslover give one to @mashe then mashe give one to @lgbasallote

mathslover (mathslover):

i gave

OpenStudy (lgbasallote):

that's an unusual way to solve @mathslover nice ;)

mathslover (mathslover):

thanks @lgbasallote

OpenStudy (anonymous):

i gave my medal too early...